OpenStudy (anonymous):

A standard deck of playing cards contains 52 cards, equally divided among four suits (hearts, diamonds, clubs, and spades). Each suit has the cards 2 through 10, as well as a jack, a queen, a king, and an ace. If the 3 of spades is drawn from a standard deck and is not replaced, what is the probability that the next card drawn is a spade OR a king?

OpenStudy (sidsiddhartha):

there are total 52 cards now if u removed the 3 of spades Our deck of cards now has 51 cards with 12 spades now chances of drawing a spade=12/51 and drawing a king is=4/51 again one of the kings is a sapde this probability is=1/51 which we cannot include so ur required prob=12/51+4/51-1/51=15/51